OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 4(sqrt3pi)

OpenStudy (saifoo.khan):

find the Radius first.

OpenStudy (anonymous):

Use the equation for circumference C = 2pi*r. Solve for r and then use that value in the equation for area A = pi*r^2

OpenStudy (saifoo.khan):

C = pi d 4(sqrt3pi) -------- = d pi

OpenStudy (saifoo.khan):

d = 12.28 r = 6.14units

OpenStudy (saifoo.khan):

Area = pi r^2 = pi (6.12)^2 =118.44unitssq

OpenStudy (anonymous):

@saifoo "exact area"